OpenStudy (anonymous):

I will give you a medal if can help me Solve the system using the elimination method. 1/2x+3y=9 1/3x+y=4 Show work PLEASE :)

OpenStudy (curry):

1/2x + 3y = 9 3(1/3x + y = 4)

OpenStudy (curry):

1/2x + 3y+ 9 x + 3y= 12

OpenStudy (curry):

-1/2x = -3

OpenStudy (curry):

1.5 = x

OpenStudy (curry):

dont forget medal

OpenStudy (anonymous):

Thanks i got you :)